A petroleum company has two different sources of crude oil. the first source provides crude oil that is 25% hydrocarbons and the second one provides crude oil that is 50% hydrocarbons. In order to obtain 60 gallons of crude oil that is 30% hydrocarbons how many gallons of crude oil must be used from each of the two sources
first source =
Second source =

Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

A petroleum company has two different sources of crude oil. the first source provides crude oil that is 25% hydrocarbons and the second one provides crude oil that is 50% hydrocarbons. In order to obtain 60 gallons of crude oil that is 30% hydrocarbons how many gallons of crude oil must be used from each of the two sources
-----
Using 2 equations::
Quantity:: t + f = 60 gallons
Oil equations: 0.25t + 0.50f = 0.30*60 oil
--------
Modify for eliminations::
25t + 25f = 25*60
25t + 50f = 30*60
---------------------
Subtract and solve for "f"::
25f = 5*60
f = 12 gallons (amt. of 50% hydro needed)
t = 60-12 = 48 gallons (amt. of 25% hydro needed)
------------
Cheers,
Stan H.
----------------





first source =
Second source =

Answer by addingup(3677)

Let's call the first source x and the second source y:
x+y= 60 subtract x from both sides:
y= 60-x
----------------------
.25x+.50y= .30(60)
.25x+.50(60-x)= .30(60)
.25x+30-.50x= 18 Now subtract x on left and subtract 30 on both sides:
-.25x= -12 Divide both sides by -.25, and remember that -/- = +
x= 48 You need 48 gallons of .25 and 60-48= 12 gallon of .5